# Environmental Breakdown of Topological Interface States in Armchair Graphene Nanoribbon Heterostructures: Is it true?

### v1

@@ Abstract
Conclusion–In this work, we theoretically investigated the transport properties of 9-7-9 AGNRH laterally embedded in boron nitride (BN) sheets. Two sandwich configurations were considered, $n$-BNNR/AGNRH/$n$-BNNR and $n$-BNNR/AGNRH/$n$-NBNR, corresponding to same-topology and reverse-topology environments, respectively. These distinct topological scenarios lead to fundamentally different behaviors of the end states (ESs) and interface states (IFs) in the AGNRH.
For same-topology BN environments, interfacial interactions between BN and C atoms strongly break the chirality of the IFs, resulting in the destruction of their topological character. In contrast, when the surrounding BN nanoribbons realize a reverse-topology configuration, the IF chirality is only weakly perturbed, allowing the IFs in $n$-BNNR/9-7-9-AGNRH/$n$-NBNR to retain the functionality of TDQDs, with energy levels that remain well isolated from the bulk states. Within a bulk boundary perturbation framework, we further demonstrate that these IFs are robust not only under vacuum boundary conditions but also against strong lateral interface interactions.
@@
Our results establish topology-engineered BN environments as an effective route to stabilizing and even enhancing topological interface states in AGNRHs. Such symmetry-protected IFs provide a promising platform for realizing robust quantum functionalities, including high-temperature quantum-dot and qubit-like applications, in graphene-based nanoelectronic devices.

### v2

@@ IV Conclusion
In this work, we theoretically investigated the electronic and transport properties of 9-7-9 and 15-13-15 AGNRHs laterally embedded in boron nitride (BN) sheets. Two sandwich configurations were considered: $n$-BNNR/AGNRH/$n$-BNNR and $n$-BNNR/AGNRH/$n$-NBNR, corresponding to same-topology and reverse-topology environments, respectively. These distinct topological configurations give rise to fundamentally different behaviors of the ESs and IFs in the AGNRH.
Within a bulk-boundary perturbation framework, we demonstrate that the topology of the surrounding BN nanoribbons plays a decisive role in determining the stability of the IFs. For same-topology BN environments, interfacial coupling between BN and C atoms induces asymmetric self-energies on the two sublattices. This effect strongly lifts the energy level of the IF state with A-chirality relative to that with B-chirality, breaking the chiral symmetry of the IF subspace and destroying their functionality as TDQDs. As a consequence, the IF levels hybridize with bulk states and lose their spectral isolation.
In contrast, when the surrounding BN nanoribbons realize a reverse-topology configuration, the global inversion symmetry is restored through mirror-inverted interfaces. In this case, the self-energy corrections from the BN environment largely cancel, leaving the IF energy levels only weakly perturbed. The IFs in $n$-BNNR/AGNRH/$n$-NBNR therefore remain well separated from the bulk continuum and retain their TDQD functionality, as confirmed by robust resonant transport reaching the conductance quantum.
@@
Our results establish topology-engineered BN environments as an effective route for stabilizing and even enhancing topological interface states in AGNRHs. The demonstrated symmetry-preserved IFs provide a promising platform for realizing robust quantum functionalities in graphene-based nanoelectronic devices, including high-temperature quantum-dot operation[ChouSY, 39,ChouSY1, 40] and qubit-like applications[LossD, 41Davidkuo5, 43].
Data availability
The data presented in this study are available upon reasonable request.
Conflicts of interest
There are no conflicts to declare
Appendix A Wider width AGNRHs embedded into boron nitride sheet
In the main text, the electronic properties of two sandwich structures, $n$-BNNR/9-7-9-AGNRH/$n$-BNNR and $n$-BNNR/9-7-9-AGNRH/$n$-NBNR, were calculated and analyzed. The former corresponds to a same-topology configuration in which the lower and upper BNNRs share identical topology, whereas the latter represents a reverse-topology configuration. Although only 9-7-9-AGNRH segments have been experimentally synthesized [DJRizzo, 10], in this appendix we investigate the environmental effects on the transport properties of 15-13-15-AGNRHs in order to examine the size dependence of AGNRH responses to BN embedding.
Figure A.1 presents the calculated energy levels of 4-BNNR/15-13-15-AGNRH/4-BNNR and 4-BNNR/15-13-15-AGNRH/4-NBNR segments as functions of the interfacial hopping strength $t_{in}$. The overall trends for the 15-13-15-AGNRH in Fig. A.1 closely resemble those of the 9-7-9 AGNRH shown in Figs. 3(b) and 3(d). However, the bulk band gap of 15-13-15-AGNRH is smaller than that of 9-7-9-AGNRH in both the same-topology and reverse-topology configurations.
In 15-13-15-AGNRHs, six in-gap energy levels appear, consisting of four ESs and two IFs. This follows from the fact that each 15-AGNR segment hosts two left ESs with A-sublattice chirality and two right ESs with B-sublattice chirality. The number of ESs for an $N$-AGNR segment with $R_{1}$ unit cells is given by $N_{es}=(N-3)/6$ and $N_{es}=(N-1)/6$ for $3p$-width and $(3p+1)$-width AGNRs, respectively. For AGNR segments with $R_{2}$ unit cells, the number of ESs becomes $N_{es}+1$. Consequently, the central 13-AGNR segment possesses three ESs at each terminus. The number of IFs at each junction is determined by $N_{IF}=|N_{es,cen}-N_{es,out}|$, where $N_{es,cen}$ and $N_{es,out}$ denote the numbers of ESs in the central and outer AGNR segments, respectively [DavidKuo, 12]. At the 15-13 junction, one IF state with A-chirality emerges, whereas at the 13-15 junction one IF state with B-chirality appears.
In the same topology configuration shown in Fig. A. 1(a), the energy-level separation $\Delta_{IF}$ between $\Sigma_{IF,C}$ and $\Sigma_{IF,V}$ for $t_{in}=0$ originates from the wave-function overlap between the A-chirality and B-chirality IFs . We note that $\Delta_{IF}$ in 15-13-15-AGNRH segments is substantially larger than that in 9-7-9-AGNRH segments, reflecting stronger intrinsic coupling between the two IFs. When $t_{in}\leq 0.1~t$, $\Delta_{IF}$ remains nearly unchanged, whereas for $t_{in}>0.2~t$, $\Delta_{IF}$ exhibits an approximate $t_{in}^{2}$ dependence. This behavior is attributed to the gradual decoupling between the A-chirality and B-chirality IFs as their spatial decay lengths are modified (see Fig. A.2(c)).
In the reverse-topology configuration shown in Fig. A.1(b), $\Delta_{IF}=2t_{eff,LR}$ remains nearly constant for $t_{in}\leq 0.3~t$. At $t_{in}=t$, we obtain $E_{g}=0.9$ eV and $\Delta_{IF}=0.362$ eV for the $15_{6}$-$13_{8}$-$15_{6}$-AGNRH segment, compared with $E_{g}=0.978$ eV and $\Delta_{IF}=0.264$ eV for the $9_{6}$-$7_{8}$-$9_{6}$-AGNRH segment. According to Eq. (4), the variation of $\Delta_{IF}$ with respect to $t_{in}$ is governed by the effective inter-IF coupling $t_{eff,LR}$. The results of Fig. A.1(b) therefore indicate that the coherent length of IFs in 15-13-15-AGNRH segments is longer than that in 9-7-9-AGNRH segments, even in the presence of BN embedding.
[…]
In Fig. A.1(a), the energy levels of the topological states are shifted due to the polarization induced by the BN sheets. To further elucidate the polarization effects arising from BNNRs in the same-topology scenario, we plot the charge densities of $\Sigma_{IF,C}$ for the $15_{6}$-$13_{8}$-$15_{6}$ AGNRH at different interfacial hopping strengths $t_{in}$ in Fig. A.2. Figures A.2(a)-2(f) correspond to $t_{in}=0$, $t_{in}=0.2~t$, $t_{in}=0.4~t$, $t_{in}=0.6~t$, $t_{in}=0.8~t$, and $t_{in}=t$, respectively. For $t_{in}=0$ and $t_{in}=0.2~t$, the charge densities of $\Sigma_{IF,C}$ are mostly confined within the graphene ribbon. However, even at $t_{in}=0.2~t$, the A-sublattice sites exhibit a noticeable enhancement of charge density, indicating the onset of BN-induced polarization. When $t_{in}=0.4~t$, the charge density becomes predominantly localized on the A-sublattice sites, which reflects the significant self-energy contributions $\Sigma_{Ni}$ and $\Sigma_{Bo}$ in Eq. (5).
As $t_{in}$ increases further to 0.6t0.6~t, 0.8t0.8~t, and $t$, the charge density of $\Sigma_{IF,C}$ extends to the B-sublattice sites at the BN edges (corresponding to nitride atom positions). Simultaneously, the B-sublattice sites at rows $i=5$ and $i=19$ in the left 15-AGNR segment also display enhanced charge density. This indicates a strong coupling between the A- and B-sublattice sites within the left 15-AGNR segment, resulting in a wave-function delocalization of $\Sigma_{IF,C}$ across the ribbon and into the BN edges.
[…]
To examine the transport properties of the 4-BNNR/15-13-15-AGNRH/4-BNNR segment, we calculate the transmission coefficient ${\cal T}_{GNR}(\varepsilon)$ as a function of applied voltage, which generates a uniform electric field along the negative $y$-direction, as shown in Fig. A.3 for interfacial hopping $t_{in}=0.8~t$ [DingY, 27GSSeal, 29]. The potential induced by the electric field is incorporated into the system Hamiltonian (Eq. (2)) as $U_{y}=-eF_{y}y$, where $F_{y}=V_{y}/L_{a}$, $V_{y}$ is the applied bias, and $L_{a}$ is the length of the AGNRH segment[DavidKuo3, 36,DavidKuo4, 37].
In Fig. A. 3(a) at $V_{y}=0$, two small transmission peaks labeled $\Sigma_{IF,C}$ and $\Sigma_{IF,V}$ appear within the bulk gap. Charge transport through these IF levels can be effectively modeled as a single quantum dot (QD) with one energy level and asymmetric tunneling rates $\Gamma_{e,L}\neq\Gamma_{e,R}$ [Davidkuo5, 43]. This demonstrates that the IFs in 4-BNNR/15-13-15-AGNRH/4-BNNR behave as a single TQD at zero bias. Increasing the tunneling rates between the zigzag edge structures of the AGNRH and the electrodes, from $\Gamma_{t}=0.56~\text{eV}$ to $2.7~\text{eV}$, enhances the peak amplitudes, although ${\cal T}_{GNR}$ remains relatively small due to the asymmetry of $\Gamma_{e,L}$ and $\Gamma_{e,R}$, as suggested by the charge density distribution in Fig. A.2(e). As illustrated in Figs. A.3(b)-(d), the separation between the $\Sigma_{IF,C}$ and $\Sigma_{IF,V}$ peaks decreases as $V_{y}$ increases. Notably, ${\cal T}_{GNR}$ reaches unity at $V_{y}=1.215$ V with $\Gamma_{t}=0.56~\text{eV}$ in Fig. A.3(d). These results indicate that the 4-BNNR/15-13-15-AGNRH/4-BNNR segment undergoes a transition from a single TQD to a TDQD under electric-field modulation. Finally, we would like to point out that our conclusions remain valid even in the case of $\Delta_{B}\neq-\Delta_{N}$. For example, when the energy level of boron atoms is $\Delta_{B}=2.329~eV$ and the energy level of nitride atoms is $\Delta_{N}=-2.499~eV$, the band inversion symmetry shown in Fig. A.1 is slightly lifted. However, the transmission coefficient still exhibits the resonant tunneling characteristic associated with topological interface states in the 4-BNNR/AGNRH/4-NBNR configuration.
[…]